OpenStudy (mydilemma):

Can someone help me with this? Find the distance between points m(6, 16) and z(-1, 14) to the nearest tenth

pooja195 (pooja195):

\[\huge~\rm~\sqrt{(x_2-x_1)^2+(y_2-y_1 )^2}\]

pooja195 (pooja195):

This is the distance formula can you plug in the numbers? :)

OpenStudy (mydilemma):

(6^2 - -1^1)^2 + (16^2 - 14^1)^2 (36 - 1)^2 + (256 - 14)^2 35^2 + 242^2 1,225 + 58,564 = 59, 789 I think I did it wrong...

pooja195 (pooja195):

Why did you sqrt everything? :O

OpenStudy (mydilemma):

Well after I put the numbers in the formula I tried solving it... what am I doing wrong?

OpenStudy (abb0t):

What are the choices?

OpenStudy (mydilemma):

There is no choices

pooja195 (pooja195):

You dont need to sqrt everything

OpenStudy (abb0t):

Oh, in that case, it's 7.

OpenStudy (mydilemma):

Okay but did i set the formula right?

OpenStudy (mydilemma):

@abb0t how did you get that?

pooja195 (pooja195):

sqrt((1-6)^2 + (14-16)^2)

OpenStudy (abb0t):

Not quite: \(\sf \color{red}{\sqrt{(-1-6)^2+(14-16)^2}}\) which is approximately 7.28, but rounding to the nearest tenth is 7.

OpenStudy (abb0t):

\(\huge \sf \color{red}{7}\) :)

pooja195 (pooja195):

Thanks for the answer abb0t we didnt need it though

OpenStudy (abb0t):

If you have any more questions, @MyDilemma if you have anymore, ask me :)

OpenStudy (mydilemma):

Thank you both for helping me, and helping me understand (::

pooja195 (pooja195):

.-.